Records over the past year show that 1 out of 315 loans made by Mammon Bank have defaulted. Find the probability that 1 or more out of 290 loans will default.

Hint: Is it appropriate to use the Poisson approximation to the binomial distribution? (Round λ to 1 decimal place. Use 4 decimal places for your answer.)
Answer #1

Yes, it is appropriate to use the Poisson approximation to the binomial distribution as n is large and p is small, giving np < 10, the necessary condition required for this approximation to work.

The Poisson parameter then becomes: X= 290 * 1/315 = 0.920634921 0.9 with

e-0.9 *0.94 P(X = z) =

We need to compute

Pr(x > 1).

Therefore, the following is obtained:

Prix > 1) = 1 - Prix <1) = 1 - Pr(X = 0)

→ Pr(x > 1) = 1-e-0.9 = 1-0.4066 = 0.5934

Hence the required probability is 0.5934

Please upvote. Let me know in the comments if anything is unclear. I will reply ASAP!